Find the vector equation of the plane determined by the points (-1,-2,1),(1,0,1) and (1,-1,1). Also check whether (1/2,1/2,1/2) lies on it.


1
Expert's answer
2021-05-26T16:36:56-0400

Solution:-

points  (-1,-2,1),(1,0,1) and (1,-1,1).

we will use this formula

"\begin{vmatrix}
 x - x_A & y - y_A&z - z_A \\
x_B - x_A & y_B - y_A & z_B - z_A\\ 
x_C - x_A &y_C - y_A & z_C - z_A
\end{vmatrix}=0"


on putting the values


"\begin{vmatrix}
x - (-1) & y - (-2)&z - 1 \\
1 - (-1) & 0 - (-2) &1 - 1\\ 
1 - (-1) &(-1) - (-2) & 1 - 1
\end{vmatrix}=0"

"\implies"

"[x - (-1)][2\u00b70-0\u00b71]-[y - (-2)][2\u00b70-0\u00b72]+[z - 1][2\u00b71-2\u00b72]=0"

"\implies"

"\boxed{- 2z + 2 = 0}"

this is the required equation

Now we check the points "(\frac{1}{2},\frac{1}{2},\frac{1}{2})"

"=-2\times\frac{1}{2}+2\\
=-1+2\\
=1 \ne0"

"\implies" points does not lie on plane.